# CSI-GPT: Integrating Generative Pre-Trained Transformer with Federated-Tuning to Acquire Downlink Massive MIMO Channels

### v1

@@ Abstract
In massive multiple-input multiple-output (MIMO) systems, how to reliably acquire downlink channel state information (CSI) with low overhead is challenging. In this work, by integrating the generative pre-trained Transformer (GPT) with federated-tuning, we propose a CSI-GPT approach to realize efficient downlink CSI acquisition. Specifically, we first propose a Swin Transformer-based channel acquisition network (SWTCAN) to acquire downlink CSI, where pilot signals, downlink channel estimation, and uplink CSI feedback are jointly designed. Furthermore, to solve the problem of insufficient training data, we propose a variational auto-encoder-based channel sample generator (VAE-CSG), which can generate sufficient CSI samples based on a limited number of high-quality CSI data obtained from the current cell. The CSI dataset generated from VAE-CSG will be used for pre-training SWTCAN. To fine-tune the pre-trained SWTCAN for improved performance, we propose an online federated-tuning method, where only a small amount of SWTCAN parameters are unfrozen and updated using over-the-air computation, avoiding the high communication overhead caused by aggregating the complete CSI samples from user equipment (UEs) to the BS for centralized fine-tuning. Simulation results verify the advantages of the proposed SWTCAN and the communication efficiency of the proposed federated-tuning method.
@@ I Introduction
In massive multiple-input multiple-output (MIMO) systems, accurate downlink channel state information (CSI) is crucial for beamforming and resource allocation. However, in frequency division duplexing (FDD) systems, accurate estimation and feedback of downlink CSI with low pilot/feedback overhead is challenging, due to the high-dimensional CSI caused by the large number of antennas at the base station (BS) and the non-reciprocity between uplink and downlink channels [1].
As the channel gains associated with different antennas are correlated, the massive MIMO CSI matrix has the inherent redundancy, which can be exploited to reduce the CSI acquisition overhead [1]. Due to its powerful capabilities of feature perception and extraction, deep learning (DL) has been widely used to process CSI in massive MIMO systems for various tasks [2]. The authors of [3] proposed a DL-based joint pilot design and channel estimation (CE) scheme, where the fully connected (FC) layer and the convolutional neural network (CNN) are used to design the pilot signal and estimate the CSI, respectively. The authors of [4] made improvements to [3] by designing pilot signals on different subcarriers differently to further reduce the pilot overhead. As for CSI feedback, the seminal work [5] proposed an autoencoder (AE)-based end-to-end (E2E) optimization framework, and the authors of [6] improved CSI feedback performance by introducing receptive fields of different sizes for CSI feature extraction.
@@
The current DL models [3, 4, 5, 7, 6, 8, 9] need a large amount of high-quality CSI samples for training, which can be obtained from actual measurements or generated from classical channel models, e.g., COST 2100 and clustered delay line (CDL) channel [11]. However, it is either difficult or communication-inefficient to obtain a large number of actual CSI samples in various practical scenarios [12], and it may degrade the NN performance if the features of training dataset and test dataset are not consistent. To overcome this issue, the generative adversarial network (GAN) is employed in [12] to generate CSI training datasets based on only a small amount of CSI measurements. Another promising solution [13] is to use federated learning (FL) to avoid collecting high-dimensional actual CSI samples, thereby reducing the communication overhead considerably. By exploiting gradient compression and over-the-air computation (AirComp), the authors of [14] proposed a communication-efficient FL framework for image classification tasks. To tune a large NN model, e.g., Transformer, in a communication-efficient manner, the authors of [15] utilized FL to fine-tune part of the parameters of a pre-trained Transformer model. The authors of [16, 17, 18] showed the potential of FL-based CE and/or CSI feedback for massive MIMO systems. They also showed that users’ data privacy can be protected by using FL. However, whether FL is more communication-efficient than conventional centralized learning (CL) that requires the feedback of CSI samples from user equipment (UEs) to BS remains unexplored.
@@ II-B Pilot Design and CSI Acquisition
The structure of the proposed SWTCAN is shown in Fig. 1. We use a linear layer without bias to model the pilot design. The received pilot signal is passed through a linear layer, and its dimension is restored to be the same as the channel. The high-dimensional features extracted by Swin Transformer blocks [19] under windows of different sizes are compressed into codewords after the linear layer and sigmoid activation function. The compressed codeword is then quantized into $B$ bits vector $\mathbf{q}$ by a quantization layer. The entire CSI compressor can be expressed as
@@ III-B2 AirComp for Efficient Federated-Tuning
To accelerate federated-tuning convergence and minimize the uplink communication rounds, we employ the federated AMSGrad with max stabilization (FedAMS) [21] in conjunction with AirComp. In the $t$-th communication round, $1\leq t\leq T$, the BS broadcasts the learnable model parameter $\tilde{\bm{\theta}}_{t}$ to all UEs (line 5 in Algorithm 1). Due to the heterogeneous user availability, only a fraction of UEs, denoted as $\mathcal{S}$, participates in the $t$-th communication round. The $u$-th UE, $\forall u\!\in\!\mathcal{S}$, initializes its parameters as $\tilde{\bm{\theta}}^{u,0}_{t}\!=\!\tilde{\bm{\theta}}_{t}$, and then minimizes its local loss function by conducting $K$ local training epochs with the local learning rate $\eta_{l}$ through its local dataset (lines 69). In the $k$-th local training epoch, $1\leq k\leq K$, learnable parameters $\tilde{\bm{\theta}}^{u,k}_{t}$ can be updated by stochastic gradient descent (SGD), which is expressed as
$\tilde{\bm{\theta}}^{u,k}_{t}=\tilde{\bm{\theta}}^{u,k-1}_{t}-\eta_{l}\nabla L_{1}^{u}(\tilde{\bm{\theta}}^{u,k-1}_{t}),$ (9)
where $\nabla$ denotes the gradient operator. After $K$ local training epochs, the $u$-th UE sends the model difference $\tilde{\bm{\theta}}^{u,K}_{t}\!-\!\tilde{\bm{\theta}}^{u,0}_{t}$ to the BS, and the BS receives the sum of the local model differences of multiple devices based on AirComp (line 10 in Algorithm 1), which can be expressed as
$\bm{\delta}_{t}=\frac{1}{|\mathcal{S}|_{c}}\sum\nolimits_{u\in\mathcal{S}}\big{(}\tilde{\bm{\theta}}^{u,K}_{t}-\tilde{\bm{\theta}}^{u,0}_{t}\big{)}+\mathbf{n}^{+},$ (10)
where $\mathbf{n}^{+}\!\in\!\mathbb{R}^{d}$ is the noise in the uplink aggregation process. According to [21], the BS updates the learnable parameters $\tilde{\bm{\theta}}_{t+1}$ for the ($t$+1)-th round (line 11 in Algorithm 1) as
$\tilde{\bm{\theta}}_{t+1}=\tilde{\bm{\theta}}_{t}+\eta\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}},$ (11)
@@ IV-A Performance of Proposed SWTCAN
We compare the proposed SWTCAN with the following baselines. Baseline 1: The Transformer-based network for CE and CSI feedback [9], denoted as ‘Transformer CE + Transformer Feedback’. Baseline 2: The multiple-measurement-vectors (MMV)-learned approximate message passing (LAMP) algorithm [22] for CE and the bit-level CsiNet scheme with an attention mechanism [23] or a Transformer-based network for CSI feedback, denoted as “MMV-LAMP CE + CsiNet/Transformer Feedback’. Baseline 3: The perfect downlink channel estimate and the bit-level CsiNet scheme with an attention mechanism [23] or a Transformer-based network for CSI feedback, denoted as ‘Perfect CE + CsiNet/Transformer Feedback’. Baseline 4: The MMV-LAMP algorithm or a Transformer-based network for CE and the perfect CSI feedback to the BS, denoted as ‘MMV-LAMP/Transformer CE + Perfect Feedback’.
Fig. 2 shows the NMSE performance achieved by different schemes versus the feedback overhead B. Both the training and test datasets are CDL-B channels. Fig. 2a indicates that at a compression ratio $\rho\!=\!\frac{N_{\rm{BS}}}{M}\!=\!8$, the main factor influencing the performance is the CSI feedback scheme. Our SWTCAN demonstrates excellent performance across all feedback bit numbers B, outperforming the baseline schemes and approaching the performance of perfect CSI feedback at $B\!=\!2048$. Fig. 2b shows that at a higher $\rho\!=\!16$, the CE algorithm also affects the final performance. Simulations on CDL-A and CDL-C channels, not showing due to space limits, also verify the same advantages of our proposed SWTCAN.
@@ IV-B Performance of Proposed Generative Pre-Training
We evaluate the performance of generative pre-training on SWTCAN at $\rho\!=\!8$ and $B\!=\!512$, 1024 and 2048 bits with the value of $l$ in (III-A) set to 0.00025. We assume that the BS has 6000 CDL-A CSI samples, which are obtained from the channel model generator. But the true CSI distribution in the BS’s cell follows a different CDL-B distribution. The BS only has 120 high-quality CDL-B samples, which are obtained from the uplink CE. We compare the proposed scheme and three benchmark schemes for ablation study. Scheme A: We pre-train SWTCAN with 6000 CDL-A samples. Scheme B: We pre-train SWTCAN with 120 CDL-B samples. Scheme C: We train VAE-CSG with 120 CDL-B samples, and then use it to generate 6000 CSI samples to pre-train SWTCAN. By contrast, in the Proposed scheme, we pre-train VAE-CSG with 6000 CDL-A samples, fine-tune it with 120 CDL-B samples, and finally generate 6000 CSI samples to pre-train SWTCAN. These four pre-training schemes are tested on 2000 CDL-B samples, and the pre-training test NMSEs are compared in Table I. It can be seen that the proposed VAE-CSG outperforms the other three benchmark schemes, demonstrating its effectiveness. Based on the proposed pre-training strategy, the performance of SWTCAN at $\rho\!=\!8$ and $B\!=\!2048$ bits reaches the NMSE of -9.3678 dB, which is still a bit short of -15.7 dB shown in Fig. 2a. Therefore, we use federated-tuning to further improve the performance.
@@
Number of feedback bits Scheme A Scheme B Scheme C Proposed
B=512 0.2009 -2.8306 -4.0628 -5.0361
B=1024 0.2491 -3.7469 -5.5036 -7.3138
B=2048 -0.1258 -3.9417 -6.6860 -9.3678
@@ IV-C Performance of Proposed Federated-Tuning Method
In this ablation study, there are $U\!=\!600$ UEs, each having $N_{\rm FL}^{s}=10$ actual CSI samples. During each communication round of federated-tuning, 10% of UEs, i.e., 60 UEs, are involved, and each UE conducts $K\!=\!2$ local training epochs with the local training rate $\eta_{l}\!=\!0.001$ to facilitate online fine-tuning of SWTCAN at $\rho\!=\!8$ and $B\!=\!2048$ bits. The hyperparameters in (11) are set to $\eta\!=\!1$, $\beta_{1}\!=\!0.9$, and $\beta_{2}\!=\!0.99$. We freeze the parameters of the SWTCAN except for the last two layers in the decoder. SNR is set to 20 dB for both downlink CE and uplink AirComp. For federated-tuning, computation (updating the trainable parameters $\tilde{\bm{\theta}}$) is done in the UE, and the model updates of multiple UEs are transmitted using AirComp [^2] Note that the uplink communication overhead can be further reduced using gradient compression methods [14]. Due to space limitation, this point is not discussed in this paper, which will be investigated in future.. For CL, the BS collects CSI samples via orthogonal transmission, and then updates the whole model parameters $\bm{\theta}$. Note that $|\tilde{\bm{\theta}}|_{c}\!=\!$ 3 617 280, $|\bm{\theta}|_{c}\!=\!$ 32 623 524, hence $|\tilde{\bm{\theta}}|_{c}\approx 0.11|\bm{\theta}|_{c}$.
@@
[b] Items Schemes Federated-tuning (one global epoch) CL (one CSI sample) Communication overhead (×$\times 10^{5}$ real numbers) $|\tilde{\bm{\theta}}|\approx 36$ $2PN_{\mathrm{BS}}\approx 1.3$ Computation cost (×$\times 10^{9}$ FLOPs) 33footnotemark: 3 $\zeta_{\mathrm{UE}}\approx 2.6$ $\zeta_{\mathrm{BS}}\approx 3.4$ Computational speed (FLOPs/s) $\kappa_{\mathrm{UE}}$ $\kappa_{\mathrm{BS}}=\gamma\kappa_{\mathrm{UE}}$
@@
Table II characterizes the communication overhead, computation cost and computation speed of the federated-tuning (one global epoch) and the CL scheme (one CSI sample). For a fair comparison, later we will compare the NMSE performance of the two schemes, given the same communication resources and the same computation time. Here, we first calculate how many CSI samples the BS can collect in CL, given a fixed communication resource. Specifically, consider $T_{0}$ global epochs of federated-tuning, within which period the communication overhead is $T_{0}|\tilde{\bm{\theta}}|_{c}$. Using the same communication resources, the BS in CL can collect $N_{\rm CL}^{s}\!=\!\frac{T_{0}|\tilde{\bm{\theta}}|_{c}}{2PN_{\mathrm{BS}}}$ CSI samples for central training. Furthermore, we calculate how many training epochs CL can have, given a fixed computation time. In particular, given the number of global epochs $T_{0}$, the computation cost $\zeta_{\mathrm{UE}}$, the number of local CSI samples $N_{\rm FL}^{s}$, the number of local training epochs $K$, and the computation speed $\kappa_{\mathrm{UE}}$, the computation time of federated-tuning can be calculated as $\tau\!=\!\frac{T_{0}N_{\rm FL}^{s}K\zeta_{\mathrm{UE}}}{\kappa_{\mathrm{UE}}}$. Within the same computation time $\tau$ in CL, given the computation cost $\zeta_{\mathrm{BS}}$, the number of CSI samples collected in the BS $N_{\rm CL}^{s}$, and the computation speed in the BS $\kappa_{\mathrm{BS}}$, we can obtain the number of training epochs of CL as $K_{\rm CL}\!=\!\frac{\tau\kappa_{\mathrm{BS}}}{N_{\rm CL}^{s}\zeta_{\mathrm{BS}}}$. Note that $\kappa_{\mathrm{BS}}\!=\!\gamma\kappa_{\mathrm{UE}}$. Although the BS is typically computationally more powerful than a UE, i.e., $\gamma>1$, our proposed federated-tuning method still shows advantages due to the collaboration of multiple UEs.

### v2

@@ Abstract
In massive multiple-input multiple-output (MIMO) systems, how to reliably acquire downlink channel state information (CSI) with low overhead is challenging. In this work, by integrating the generative pre-trained Transformer (GPT) with federated-tuning, we propose a CSI-GPT approach to realize efficient downlink CSI acquisition. Specifically, we first propose a Swin Transformer-based channel acquisition network (SWTCAN) to acquire downlink CSI, where pilot signals, downlink channel estimation, and uplink CSI feedback are jointly designed. Furthermore, to solve the problem of insufficient training data, we propose a variational auto-encoder-based channel sample generator (VAE-CSG), which can generate sufficient CSI samples based on a limited number of high-quality CSI data obtained from the current cell. The CSI dataset generated from VAE-CSG will be used for pre-training SWTCAN. To fine-tune the pre-trained SWTCAN for improved performance, we propose an online federated-tuning method, where only a small amount of SWTCAN parameters are unfrozen and updated using over-the-air computation, avoiding the high communication overhead caused by aggregating the complete CSI samples from user equipment (UEs) to the BS for centralized fine-tuning. Simulation results verify the advantages of the proposed SWTCAN and the communication efficiency of the proposed federated-tuning method. Our code is publicly available at https://github.com/BIT-ZY/CSI-GPT
@@ I Introduction
In massive multiple-input multiple-output (MIMO) systems, accurate downlink channel state information (CSI) is crucial for beamforming and resource allocation. However, in frequency division duplexing (FDD) systems, accurate estimation and feedback of downlink CSI with low pilot/feedback overhead is challenging, due to the high-dimensional CSI caused by the large number of antennas at the base station (BS) and the non-reciprocity between uplink and downlink channels [1, 2, 3, 4].
As the channel gains associated with different antennas are correlated, the massive MIMO CSI matrix has the inherent redundancy, which can be exploited to reduce the CSI acquisition overhead [5, 6, 7, 8]. Due to its powerful capabilities of feature perception and extraction, deep learning (DL) has been widely used to process CSI in massive MIMO systems for various tasks. The authors of [9] proposed a DL-based joint pilot design and channel estimation scheme, where the fully connected (FC) layer and the convolutional neural network (CNN) are used to design the pilot signal and estimate the CSI, respectively. The authors of [10] made improvements to [9] by designing pilot signals on different subcarriers differently to further reduce the pilot overhead. As for CSI feedback, the seminal work [11] proposed an autoencoder (AE)-based end-to-end (E2E) optimization framework, and the authors of [12] improved CSI feedback performance by introducing receptive fields of different sizes for CSI feature extraction.
To improve the deployability of CSI feedback, the authors in [13] proposed a scheme in which the length of the feedback codeword is variable with the sparsity of the channel. The authors in [14] proposed a scheme based on knowledge distillation. Both schemes largely reduce the complexity of the network that needs to be deployed. In addition, the latest information about the application of DL in CSI feedback and the comparison can be obtained from [15].
@@
The current DL models [9, 10, 11, 16, 12, 17, 18] need a large amount of high-quality CSI samples for training, which can be obtained from actual measurements or generated from classical channel models, e.g., COST 2100 and clustered delay line (CDL) channel [20]. However, it is either difficult or communication-inefficient to obtain a large number of actual CSI samples in various practical scenarios [21], and it may degrade the NN performance if the features of training dataset and test dataset are not consistent. To overcome this issue, the generative adversarial network (GAN) is employed in [21] to generate CSI training datasets based on only a small amount of CSI measurements. Another promising solution [22] is to use federated learning (FL) to avoid collecting high-dimensional actual CSI samples, thereby reducing the communication overhead considerably. By exploiting gradient compression and over-the-air computation (AirComp), the authors of [23] proposed a communication-efficient FL framework for image classification tasks. The authors further proposed a massive digital AirComp scheme that are compatible with the current wireless networks in [24]. To tune a large NN model, e.g., Transformer, in a communication-efficient manner, the authors of [25] utilized FL to fine-tune part of the parameters of a pre-trained Transformer model. The authors of [26, 27, 28] showed the potential of FL-based CE and/or CSI feedback for massive MIMO systems. They also showed that users’ data privacy can be protected by using FL. However, whether FL is more communication-efficient than conventional centralized learning (CL) that requires the feedback of CSI samples from user equipment (UEs) to BS remains unexplored.
@@ II-B Pilot Design and CSI Acquisition
The structure of the proposed SWTCAN is shown in Fig. 1. We use a linear layer without bias to model the pilot design. The received pilot signal is passed through a linear layer, and its dimension is restored to be the same as the channel.
The core of the compressor consists of 8 Swin Transformer blocks[29]. These blocks are responsible for extracting high-dimensional features from the input signal, leveraging a 96-dimensional embedding space. Each pair of Swin Transformer blocks starts with LayerNorm, followed by shifted window multi-head self-attention (SW-MSA) in the first block and window-based multi-Head self-attention (W-MSA) in the second, both paired with multilayer perceptron (MLPs) and residual connections. This structure captures the local and global features of CSI effectively. The features passes through a linear layer with an activation function and form a codeword. The compressed codeword is then quantized into $B$ bits vector $\mathbf{q}$ by a quantization layer. The entire CSI compressor can be expressed as
@@ III-B2 AirComp for Efficient Federated-Tuning
To accelerate federated-tuning convergence and minimize the uplink communication rounds, we employ the federated AMSGrad with max stabilization (FedAMS) [31] in conjunction with AirComp.
The BS begins by initializing the SWTCAN model with pre-trained parameters. These parameters have been pre-trained using a large dataset of CSI samples generated by the VAE-CSG. In the $t$-th communication round, $1\leq t\leq T$, the BS broadcasts the learnable model parameter $\tilde{\bm{\theta}}_{t}$ to all UEs (line 5 in Algorithm 1). Due to the heterogeneous user availability, only a fraction of UEs, denoted as $\mathcal{S}$, participates in the $t$-th communication round. The $u$-th UE, $\forall u\!\in\!\mathcal{S}$, initializes its parameters as $\tilde{\bm{\theta}}^{u,0}_{t}\!=\!\tilde{\bm{\theta}}_{t}$, and then minimizes its local loss function by conducting $K$ local training epochs with the local learning rate $\eta_{l}$ through its local dataset (lines 69). In the $k$-th local training epoch, $1\leq k\leq K$, learnable parameters $\tilde{\bm{\theta}}^{u,k}_{t}$ can be updated by stochastic gradient descent (SGD), which is expressed as
$\tilde{\bm{\theta}}^{u,k}_{t}=\tilde{\bm{\theta}}^{u,k-1}_{t}-\eta_{l}\nabla L_{1}^{u}(\tilde{\bm{\theta}}^{u,k-1}_{t}),\vspace{-1mm}$ (9)
where $\nabla$ denotes the gradient operator. After $K$ local training epochs, instead of sending the entire model back to the BS, the $u$-th UE sends the model difference $\Delta\tilde{\bm{\theta}}^{u}_{t}=\tilde{\bm{\theta}}^{u,K}_{t}\!-\!\tilde{\bm{\theta}}^{u,0}_{t}$ to the BS, and the BS receives the sum of the local model differences from multiple devices based on AirComp (line 10 in Algorithm 1), which can be expressed as
$\bm{\delta}_{t}=\frac{1}{|\mathcal{S}|_{c}}\sum\nolimits_{u\in\mathcal{S}}\Delta\tilde{\bm{\theta}}^{u}_{t}+\mathbf{n}^{+},\vspace{-1mm}$ (10)
where $\mathbf{n}^{+}\!\in\!\mathbb{R}^{d}$ is the noise in the uplink aggregation process, $\bm{\delta}_{t}$ represents the noisy model difference, which is also treated as a pseudo gradient to update the model at the BS. According to [31], the BS updates the learnable parameters $\tilde{\bm{\theta}}_{t+1}$ for the ($t$+1)-th round (line 11 in Algorithm 1) as
$\tilde{\bm{\theta}}_{t+1}=\tilde{\bm{\theta}}_{t}+\eta\frac{\mathbf{m}_{t}}{\sqrt{\mathbf{v}_{t}}},\vspace{-1mm}$ (11)
@@ IV Simulation Results
(a) $\rho=2$
(b) $\rho=2$
(c) $\rho=2$
(d) $\rho=32$
(e) $\rho=32$
(f) $\rho=32$
Figure 3: NMSE performance of different schemes versus the feedback signaling overhead for CDL-A/B/C.
@@ IV-A Performance of Proposed SWTCAN
We compare the proposed SWTCAN with the following baselines. Baseline 1: The Transformer-based network for CE and CSI feedback [18], denoted as ‘Transformer CE + Transformer Feedback’. Baseline 2: The multiple-measurement-vectors (MMV)-learned approximate message passing (LAMP) algorithm [32] for CE and the bit-level CsiNet scheme with an attention mechanism [33] or a Transformer-based network for CSI feedback, denoted as ‘MMV-LAMP CE + CsiNet/Transformer Feedback’. Baseline 3: The perfect downlink channel estimate and the bit-level CsiNet scheme with an attention mechanism [33] or a Transformer-based network for CSI feedback, denoted as ‘Perfect CE + CsiNet/Transformer Feedback’. Baseline 4: The MMV-LAMP algorithm or a Transformer-based network for CE and the perfect CSI feedback to the BS, denoted as ‘MMV-LAMP/Transformer CE + Perfect Feedback’.
Fig. 2 shows the NMSE performance achieved by different schemes versus the feedback overhead B. Both the training and test datasets are CDL-B channels. Fig. 2a indicates that at a compression ratio $\rho\!=\!\frac{N_{\rm{BS}}}{M}\!=\!8$, the main factor influencing the performance is the CSI feedback scheme. Our SWTCAN demonstrates excellent performance across all feedback bit numbers B, outperforming the baseline schemes and approaching the performance of perfect CSI feedback at $B\!=\!2048$. Fig. 2b shows that at a higher $\rho\!=\!16$, the CE algorithm also affects the final performance. Simulations on CDL-A and CDL-C channels, not showing due to space limits, also verify the same advantages of our proposed SWTCAN. Thank you for your suggestion regarding the inclusion of additional experiments with different channel configurations. We agree that demonstrating the generalizability of our proposed approach across various channel models is important. While our primary simulations were conducted under the CDL-B channel, we have also performed extensive simulations using CDL-A and CDL-C channels. These additional experiments confirm the superiority of our proposed algorithm across different channel conditions. However, due to the page limit constraints (with a maximum of 7 figures and tables allowed), we were unable to include these results in the main manuscript. We have attached the relevant simulation results and analysis in this response letter for your review.
As shown in Fig. 3, we extended our simulations to include compression ratios $\rho$ of 2 and 32, in addition to the $\rho$ = 8 and $\rho$ = 16 scenarios presented in the manuscript. These extended simulations were conducted across CDL-A, CDL-B, and CDL-C channels to examine the boundary effects on overall performance, specifically regarding channel estimation and feedback algorithms. The results showed that at a compression ratio of $\rho$ = 2, the performance is primarily determined by the channel feedback algorithm. Our proposed SWTCAN algorithm approaches the performance of perfect CSI feedback when using the MMV-LAMP channel estimation algorithm. Conversely, at a compression ratio of $\rho$ = 32, channel estimation becomes the decisive factor for overall performance. Here, the SWTCAN algorithm outperforms the baseline in nearly all conditions, even with lossy channel estimation and feedback. Notably, in the CDL-B channel at $\rho$ = 32, all algorithms requiring channel estimation demonstrate weaker performance, making the Transformer-based feedback scheme with perfect channel estimation particularly effective. In contrast, the CsiNet network, due to its less robust feature extraction capability, fails to achieve satisfactory performance even under perfect channel estimation.
[…]
@@ IV-B Performance of Proposed Generative Pre-Training
We evaluate the performance of generative pre-training on SWTCAN at $\rho\!=\!8$ and $B\!=\!512$, 1024 and 2048 bits with the value of $l$ in (III-A) set to 0.00025. We assume that the BS has 6000 CDL-A CSI samples, which are obtained from the channel model generator. But the true CSI distribution in the BS’s cell follows a different CDL-B distribution. The BS only has 120 high-quality CDL-B samples, which are obtained from the uplink CE. We compare the proposed scheme and three benchmark schemes for ablation study. Scheme A: We pre-train SWTCAN with 6000 CDL-A samples. Scheme B: We pre-train SWTCAN with 120 CDL-B samples. Scheme C: We pre-train SWTCAN with 6000 CDL-A samples, fine-tune it with 120 CDL-B samples. Scheme D: We train VAE-CSG with 120 CDL-B samples, and then use it to generate 6000 CSI samples to pre-train SWTCAN. By contrast, in the Proposed scheme, we pre-train VAE-CSG with 6000 CDL-A samples, fine-tune it with 120 CDL-B samples, and finally generate 6000 CSI samples to pre-train SWTCAN. These four pre-training schemes are tested on 2000 CDL-B samples, and the pre-training test NMSEs are compared in Table I. It can be seen that the proposed VAE-CSG outperforms the other three benchmark schemes, demonstrating its effectiveness. Based on the proposed pre-training strategy, the performance of SWTCAN at $\rho\!=\!8$ and $B\!=\!2048$ bits reaches the NMSE of -9.3678 dB, which is still a bit short of -15.7 dB shown in Fig. 2a. Therefore, we use federated-tuning to further improve the performance.
@@
Number of feedback bits Scheme A Scheme B Scheme C Scheme D Proposed
B=512 0.2009 -2.8306 -3.8199 -4.0628 -5.0361
B=1024 0.2491 -3.7469 -4.8259 -5.5036 -7.3138
B=2048 -0.1258 -3.9417 -5.5449 -6.6860 -9.3678
@@ IV-C Performance of Proposed Federated-Tuning Method
In this ablation study, there are $U\!=\!600$ UEs, each having $N_{\rm FL}^{s}=10$ actual CSI samples. During each communication round of federated-tuning, 10% of UEs, i.e., 60 UEs, are involved, and each UE conducts $K\!=\!2$ local training epochs with the local training rate $\eta_{l}\!=\!0.001$ to facilitate online fine-tuning of SWTCAN at $\rho\!=\!8$ and $B\!=\!2048$ bits. The hyperparameters in (11) are set to $\eta\!=\!1$, $\beta_{1}\!=\!0.9$, and $\beta_{2}\!=\!0.99$. We freeze the parameters of the SWTCAN except for the last two layers 222 Our study showed that freezing the last two layers of all Swin Transformer Blocks (with 3,617,280 unfrozen parameters) resulted in a final performance of -10.722 dB after 25 global epochs. Freezing only the last layer (2,509,299 unfrozen parameters) led to a performance of -10.299 dB, while freezing half of the last layer (2,471,667 unfrozen parameters) achieved -10.202 dB. The chosen freezing scheme strikes a balance between communication efficiency and performance, which is verified in the simulations. in the decoder. SNR is set to 20 dB for both downlink CE and uplink AirComp. For federated-tuning, computation (updating the trainable parameters $\tilde{\bm{\theta}}$) is done in the UE, and the model updates of multiple UEs are transmitted using AirComp [^3] Note that the uplink communication overhead can be further reduced using gradient compression methods [23]. Due to space limitation, this point is not discussed in this paper, which will be investigated in future.. For CL, the BS collects CSI samples via orthogonal transmission, and then updates the whole model parameters $\bm{\theta}$. Note that $|\tilde{\bm{\theta}}|_{c}\!=\!$ 3 617 280, $|\bm{\theta}|_{c}\!=\!$ 32 623 524, hence $|\tilde{\bm{\theta}}|_{c}\approx 0.11|\bm{\theta}|_{c}$.
@@
[b] Items Schemes Federated-tuning (one global epoch) CL (one CSI sample) Communication overhead (×$\times 10^{5}$ real numbers) $|\tilde{\bm{\theta}}|_{c}\approx 36$ $2PN_{\mathrm{BS}}\approx 1.3$ Computation cost (×$\times 10^{9}$ FLOPs) 33footnotemark: 3 $\zeta_{\mathrm{UE}}\approx 2.6$ $\zeta_{\mathrm{BS}}\approx 3.4$ Computational speed (FLOPs/s) $\kappa_{\mathrm{UE}}$ $\kappa_{\mathrm{BS}}=\gamma\kappa_{\mathrm{UE}}$
@@
Table II characterizes the communication overhead, computation cost and computation speed of the federated-tuning (one global epoch) and the CL scheme (one CSI sample). For federated-tuning, communication overhead during each global epoch is denoted as $|\tilde{\bm{\theta}}|_{c}$, representing the number of trainable parameters of SWTCAN. While, the communication overhead of CL is calculated as $2PN_{\mathrm{BS}}$ for each CSI sample. The computation cost is measured in FLOPs using the torch-summary tool, where $\zeta_{\mathrm{UE}}$ and $\zeta_{\mathrm{BS}}$ represent the computation cost in federated-tuning and CL, respectively. Computation speed, in FLOPs/s, is indicated by $\kappa_{\mathrm{BS}}$ and $\kappa_{\mathrm{UE}}$ for the BS and UE, respectively, where their ratio $\gamma=\frac{\kappa_{\mathrm{BS}}}{\kappa_{\mathrm{UE}}}$ reflects the computational power difference between the BS and UEs.
For a fair comparison, later we will compare the NMSE performance of the two schemes, given the same communication resources and the same computation time. Here, we first calculate how many CSI samples the BS can collect in CL, given a fixed communication resource. Specifically, consider $T_{0}$ global epochs of federated-tuning, within which period the communication overhead is $T_{0}|\tilde{\bm{\theta}}|_{c}$. Using the same communication resources, the BS in CL can collect $N_{\rm CL}^{s}\!=\!\frac{T_{0}|\tilde{\bm{\theta}}|_{c}}{2PN_{\mathrm{BS}}}$ CSI samples for central training. Furthermore, we calculate how many training epochs CL can have, given a fixed computation time. In particular, given the number of global epochs $T_{0}$, the computation cost $\zeta_{\mathrm{UE}}$, the number of local CSI samples $N_{\rm FL}^{s}$, the number of local training epochs $K$, and the computation speed $\kappa_{\mathrm{UE}}$, the computation time of federated-tuning can be calculated as $\tau\!=\!\frac{T_{0}N_{\rm FL}^{s}K\zeta_{\mathrm{UE}}}{\kappa_{\mathrm{UE}}}$. Within the same computation time $\tau$ in CL, given the computation cost $\zeta_{\mathrm{BS}}$, the number of CSI samples collected in the BS $N_{\rm CL}^{s}$, and the computation speed in the BS $\kappa_{\mathrm{BS}}$, we can obtain the number of training epochs of CL as $K_{\rm CL}\!=\!\frac{\tau\kappa_{\mathrm{BS}}}{N_{\rm CL}^{s}\zeta_{\mathrm{BS}}}$. Although the BS is typically computationally more powerful than a UE, i.e., $\gamma>1$, our proposed federated-tuning method still shows advantages due to the collaboration of multiple UEs.